OpenStudy (anonymous):

A merchant blends tea that sells for $10.70 a pound with tea that sells for $2.20 a pound to produce 10 lbs of a mixture that sells for $8.15 a pound. The final mixture contains __________ lbs of $10.70 tea and ________ lbs of $2.20 tea.

OpenStudy (anonymous):

10.7*a + 2.2*b = 8.15*10 a + b = 10

OpenStudy (anonymous):

i dont understand.

OpenStudy (anonymous):

Let a be the amount of $10.70 tea in pounds, and b be the amount of $2.20 tea in pounds. There are a total of 10 lbs (IDK if that means pounds) so a + b = 10; The tea sells for $8.15 per pound, and the total weight of combined tea is 10 lbs so the total worth of the combined tea is 8.15*10 = 81.5; Then the total value consists of a pounds of $10.70 tea and b pounds of $2.20 tea, so you have the 1st equation.